### v1

@@ 5 Application I: Augmented Lagrangian method
In the remainder of this paper, we present several examples illustrating how the auxiliary space theory and the estimates for the Schur complements developed in Section 4 can be applied to different applications. As a first example, we consider the augmented Lagrangian method [12], which is a stationary iterative method for solving the saddle point problem (1).
@@
(15) $\begin{aligned} u^{m+1}&=(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g-B^{t}p^{m}),\\ p^{m+1}&=p^{m}-\epsilon^{-1}(g-Bu^{m+1}),\end{aligned}\quad m\geq 0.$
@@
(16) $S_{\epsilon}p=d_{\epsilon},\quad\text{where}\quad S_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}B^{t},\,d_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g)-g.$
@@
Thus we obtain the identity (cf. [18, equation (3.4)]):
$S_{\epsilon}^{-1}=S^{-1}+\epsilon^{-1}I.$
@@
Namely, the convergence rate of (15) becomes arbitrarily fast as $\epsilon$ becomes small. This result can be regarded as a refined version of [21, Lemma 2.1], with much simpler proof. We note that such arbitrarily fast convergence of the augmented Lagrangian method can be extended to a broader class of convex optimization problems; see [32, Appendix A].
@@ 6.1 Darcy flow
where $M_{W}\colon W\to W$ is the mass matrix on $W$ defined analogously to $M_{V}$. Moreover, the mass matrices $M_{V}$ and $M_{W}$ satisfy the scaling relations (cf. [37, Section 3.4])
@@
where the last inequality follows from the standard inverse inequality (see, e.g., [30, Lemma 2.2]). In conclusion, we obtain
@@ 6.2 Stokes equations
Let $\Omega$ be a bounded polygonal domain in $\mathbb{R}^{d}$ ($d=2,3$). The incompressible Stokes equations with the homogeneous Dirichlet boundary condition are written as
@@
A mixed finite element method for solving (24) is obtained by replacing $H_{0}^{1}(\Omega)^{d}$ and $L_{0}^{2}(\Omega)$ with suitable finite element spaces $V$ and $W$. Then the corresponding algebraic system becomes (1), where the matrices AA, $B$ and the vector $f$ are defined by
@@ 7.1 FETI
(29) $\min_{v=(v_{j})_{j=1}^{J}\in V}\left\{\frac{1}{2}(Sv,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=0,$
where the matrix $S$ and the vector $f$ are defined by
(30) $(Sv,w)=\sum_{j=1}^{J}\int_{\Omega_{j}}\nabla\mathcal{H}_{j}v_{j}\cdot\nabla\mathcal{H}_{j}w_{j}\,dx,\qquad(f,v)=\sum_{j=1}^{J}\int_{\Omega_{j}}f\,\mathcal{H}_{j}v_{j}\,dx,\quad v,w\in V,$
$\mathcal{H}_{j}$ denotes the discrete harmonic extension in $\Omega_{j}$ associated with $\mathcal{T}_{h}^{j}$ (see [37, 43] for details), $B$ is a full-rank matrix with entries 0 and $\pm 1$ enforcing continuity along $\Gamma$, and $(\cdot,\cdot)$ denotes the Euclidean inner product. It is straightforward to observe that (29) is equivalent to the restriction of the finite element discretization of (28) on the global mesh $\mathcal{T}_{h}$ to the interface $\Gamma$.
@@
(31) $\begin{bmatrix}S&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
Note that $S$ is semi-SPD, owing to subdomains $\Omega_{j}$ that do not intersect $\partial\Omega$. Finally, by eliminating $u$ in (31) as in Section 4, we obtain the following SPD linear system corresponding to (14):
(32) $F_{0}\lambda_{0}=d_{0},\quad\text{where}\quad F_{0}=PBS^{+}(PB)^{t}\colon W_{0}\to W_{0},\ d_{0}\in W_{0},$
@@
Now we estimate the condition number of $F_{0}$ in (32) using the proposed framework. We begin by summarizing some useful properties of the matrices $S$ and $B$. The matrix $S$ satisfies the following bounds, which combine the well-known estimate for Schur complements in the finite element method (see, e.g., [34, Lemma 1.59]) with standard scaling arguments [37, Section 3.4]:
(33) $\lambda_{\min}(S)\gtrsim\frac{h}{H},\quad\lambda_{\max}(S)\lesssim 1,$
where $\lambda_{\min}$ denotes the minimum nonzero eigenvalue. Furthermore, as shown in [10, Equation (22)], since the matrix $B$ consists only of entries 0 and $\pm 1$, an elementary calculation gives
@@
By Theorem 4.9, the minimum eigenvalue of $F_{0}$ is estimated as
@@
where in the second inequality we substituted $v=B^{t}(BB^{t})^{-1}\mu$, which satisfies $PBv=\mu$. On the other hand, the maximum eigenvalue of $F_{0}$ is estimated as
@@
$\kappa(F_{0})\lesssim\frac{H}{h}.$
This result agrees with [10, Theorem 3.2], where the condition number of $F_{0}$ was first estimated.
@@
The only technical ingredient we require is the well-known vertex–edge lemma [43, Section 4.3], stated in Lemma 7.1. In what follows, edges are understood as relatively open subsets, i.e., without their endpoints. For a subset $K\subset\partial\Omega_{j}$, the restriction $I_{K}^{0}v_{j}\in V_{j}$ of $v_{j}\in V_{j}$ is defined by
@@
    $\|I_{x}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
@@
    $\|I_{e}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
@@
To estimate the maximum eigenvalue of $L_{\mathrm{FETI}}S_{0}$, Corollary 4.12 implies that it suffices to bound $(S(P\bar{B})^{t}PBv,(P\bar{B})^{t}PBv)$ for any $v\in\mathcal{R}(S)$. Take any $v\in\mathcal{R}(S)$ and choose $\phi\in\mathcal{N}(S)$ such that $v+\phi\in W_{0}$. Then we have
@@
where the first inequality follows from the definitions of $B$ and $\bar{B}$ together with the triangle inequality, and the second inequality follows from Lemma 7.1. Here we require mild geometric assumptions on the domain decomposition $\{\Omega_{j}\}_{j=1}^{J}$, namely that the numbers of vertices and edges of the subdomains are uniformly bounded and that the number of subdomains sharing a common vertex is also uniformly bounded. Hence, by Corollary 4.12, we deduce that
$\lambda_{\max}(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
In summary, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{FETI}}F_{0}$ (cf. [37, Theorem 6.15]):
$\kappa(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
We note that the unscaled Dirichlet preconditioner, namely $(PB)S(PB)^{t}$, was first proposed in [10] and analyzed in [26]. In the unscaled case, as discussed in [26], the minimum eigenvalue admits a weaker estimate in certain scenarios.
@@ 7.2 FETI-DP
(37) $\min_{v=(v_{j})_{j=1}^{J}\in\tilde{V}}\left\{\frac{1}{2}(\tilde{S}v,v)-(f,v)\right\},\quad\text{subject to}\quad Bv=0,$
where the matrix $\tilde{S}$ and the vector $\tilde{f}$ are defined analogously to (30); see [16, 37] for details. Here $B$ is a full-rank matrix with entries 0 and $\pm 1$ that enforces continuity along the interior of subdomain edges, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
@@
(38) $\begin{bmatrix}\tilde{S}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
Different from the FETI formulation (31), in (38) the matrix $\tilde{S}$ is SPD. Therefore, by eliminating $u$ in (38), we obtain the following Schur complement system corresponding to (7):
(39) $F\lambda=d,\quad\text{where}\quad F=B\tilde{S}^{-1}B^{t},\quad d=\tilde{S}^{-1}f.$
In (37), since the matrix $B$ enforces continuity only at nodal points located in the interior of subdomain edges, each nodal point is affected by $B$ only once. Hence, we have the following identity [27, Equation (3.2)]:
@@
As FETI-DP involves functions that are continuous at subdomain corners, we require certain Poincaré-type inequalities associated with subdomain corners in order to analyze FETI-DP. In Lemma 7.2, which summarizes results from [19, Lemma 2] and [27, Lemma 5.1], $L^{2}$ and $H^{\frac{1}{2}}$ estimates are provided for functions vanishing at subdomain corners. In what follows, let $I_{H}$ denote the linear coarse interpolation operator associated with the coarse triangulation $\mathcal{T}_{H}$.
@@
Now, we estimate the extremal eigenvalues of $F$ given in (39). By Theorem 4.1, we obtain
@@
$\lambda_{\max}(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right).$
@@
$\kappa(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right),$
which agrees with the existing result in [19, Proposition 3].
@@
$L_{\mathrm{DP}}=\bar{B}\tilde{S}\bar{B}^{t},$
By Corollary 4.7, we have $\lambda_{\min}(L_{\mathrm{DP}}F)\geq 1$. Moreover, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, it follows that
$\begin{split}(\tilde{S}\bar{B}^{t}Bv,\bar{B}^{t}Bv)&=(\tilde{S}\bar{B}^{t}B(v-I_{H}v),\bar{B}^{t}B(v-I_{H}v))\\ &=\sum_{j=1}^{J}|\bar{B}^{t}B(v-I_{H}v)|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\sum_{j=1}^{J}\sum_{e\text{:~edge of }\Omega_{j}}\left|I_{e}^{0}(v_{j}-I_{H}v_{j})\right|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\left(1+\log\frac{H}{h}\right)^{2}\sum_{j=1}^{J}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &=\left(1+\log\frac{H}{h}\right)^{2}(\tilde{S}v,v),\end{split}$
@@
$\lambda_{\max}(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
Finally, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{DP}}F$ (cf. [27, Theorem 5.2]):
$\kappa(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
@@
Two closely related nonoverlapping domain decomposition methods to FETI and FETI-DP are BDD (balancing domain decomposition) [13, 23] and BDDC (balancing domain decomposition by constraints) [8, 13]. It has been shown in the literature [6, 22, 24, 25, 36] that FETI/FETI-DP with scaled Dirichlet preconditioners, and BDD/BDDC are spectrally equivalent in the sense that their spectra agree except for the eigenvalues 0 and 11. In fact, BDD and BDDC can be analyzed within our framework (Theorems 4.5 and 4.10) by essentially the same arguments used for FETI and FETI-DP. For brevity, we omit the details.

### v2

@@ 5 Application I: Augmented Lagrangian method
In the remainder of this paper, we present several examples illustrating how the auxiliary space theory and the estimates for the Schur complements developed in Section 4 can be applied to different problems. As a first example, we consider the augmented Lagrangian method [13], which is a stationary iterative method for solving the saddle point problem (1).
@@
(16) $\begin{aligned} u^{m+1}&=(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon^{-1}B^{t}g-B^{t}p^{m}),\\ p^{m+1}&=p^{m}-\epsilon^{-1}(g-Bu^{m+1}),\end{aligned}\quad m\geq 0.$
@@
(17) $S_{\epsilon}p=d_{\epsilon},\text{ where }S_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}B^{t},\,d_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon^{-1}B^{t}g)-g.$
@@
Thus we obtain the identity (cf. [20, equation (3.4)]):
(18) $S_{\epsilon}^{-1}=S^{-1}+\epsilon^{-1}I.$
@@
Namely, the convergence rate of (16) becomes arbitrarily fast as $\epsilon$ becomes small. This result can be regarded as a refined version of [23, Lemma 2.1], with much simpler proof. We note that such arbitrarily fast convergence of the augmented Lagrangian method can be extended to a broader class of convex optimization problems; see [34, Appendix A].
@@
Remark 5.1.
Alternative derivations of the identity (18) can be found in the literature. For example, the well-known Sherman–Morrison–Woodbury formula [15, equation (2.1.4)],
$(A+UCV)^{-1}=A^{-1}-A^{-1}U(C^{-1}+VA^{-1}U)^{-1}VA^{-1},$
where AA, $C$, $U$, and $V$ are conformable matrices, can be used to derive (18); see, for example, [3, Lemma 4.1] and [20, equation (3.4)].
@@ 6.1 Darcy flow
where $M_{W}\colon W\to W$ is the mass matrix on $W$ defined analogously to $M_{V}$. Moreover, the mass matrices $M_{V}$ and $M_{W}$ satisfy the scaling relations (cf. [39, Section 3.4])
@@
where the last inequality follows from the standard inverse inequality (see, e.g., [32, Lemma 2.2]). In conclusion, we obtain
@@ 6.2 Stokes equations
Let $\Omega$ be a bounded polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). The incompressible Stokes equations with the homogeneous Dirichlet boundary condition are written as
@@
A mixed finite element method for solving (26) is obtained by replacing $H_{0}^{1}(\Omega)^{d}$ and $L_{0}^{2}(\Omega)$ with suitable finite element spaces $V$ and $W$. Then the corresponding algebraic system becomes (1), where the matrices $A$, $B$ and the vector $f$ are defined by
@@ 7.1 FETI
(31) $\min_{v=(v_{j})_{j=1}^{J}\in V}\left\{\frac{1}{2}(Av,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=0,$
where the matrix $A$ and the vector $f$ are defined by
(32) $(Av,w)=\sum_{j=1}^{J}\int_{\Omega_{j}}\nabla\mathcal{H}_{j}v_{j}\cdot\nabla\mathcal{H}_{j}w_{j}\,dx,\qquad(f,v)=\sum_{j=1}^{J}\int_{\Omega_{j}}f\,\mathcal{H}_{j}v_{j}\,dx,\quad v,w\in V,$
where $\mathcal{H}_{j}$ denotes the discrete harmonic extension in $\Omega_{j}$ associated with $\mathcal{T}_{h}^{j}$ (see [39, 45] for details), $B$ is a full-rank matrix with entries 0 and $\pm 1$ enforcing continuity along $\Gamma$, and $(\cdot,\cdot)$ denotes the Euclidean inner product. It is straightforward to observe that (31) is equivalent to the restriction of the finite element discretization of (30) on the global mesh $\mathcal{T}_{h}$ to the interface $\Gamma$.
@@
(33) $\begin{bmatrix}A&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
Note that $A$ is semi-SPD, owing to subdomains $\Omega_{j}$ that do not intersect $\partial\Omega$. Finally, by eliminating $u$ in (33) as in Section 4, we obtain the following SPD linear system corresponding to (14):
(34) $S_{0}\lambda_{0}=d_{0},\quad\text{where}\quad S_{0}=PBA^{+}(PB)^{t}\colon W_{0}\to W_{0},\ d_{0}\in W_{0},$
@@
Now we estimate the condition number of $S_{0}$ in (34) using the proposed framework. We begin by summarizing some useful properties of the matrices $A$ and $B$. We have the following estimates for the extremal eigenvalues of $A$ (cf. [11, equations (29) and (30)]).
Lemma 7.1.
The matrix AA defined in (32) satisfies
(35) $\lambda_{\min}(A)\gtrsim\frac{h}{H},\quad\lambda_{\max}(A)\lesssim 1,$
where $\lambda_{\min}$ denotes the smallest nonzero eigenvalue.
Proof 7.2.
For any $v=(v_{j})_{j=1}^{J}\in\mathcal{R}(A)$, we have
[…]
which completes the proof.
Remark 7.3.
Let $\hat{V}$ be a subspace of $V$ consisting of functions that are continuous across $\Gamma$. Then the restriction $A|_{\hat{V}}$ is SPD, since continuity across $\Gamma$ eliminates the piecewise constant kernel modes associated with floating subdomains. For $A|_{\hat{V}}$, we cannot apply the Poincaré inequality on each $\partial\Omega_{j}$ as in Lemma 7.1; instead, we obtain the following weaker estimate for the minimum eigenvalue (see, e.g., [39, Lemma 4.11] and [45, Lemma 3.1]), derived using the global Poincaré inequality on $\Omega$:
$(Av,v)\gtrsim H\sum_{j=1}^{J}\|v_{j}\|_{L^{2}(\partial\Omega_{j})}^{2}\eqsim Hh(v,v),\quad v=(v_{j})_{j=1}^{J}\in\hat{V}$.
In addition, as shown in [11, equation (22)], since the matrix $B$ consists only of entries 0 and $\pm 1$, an elementary calculation gives
@@
By Theorem 4.9, the minimum eigenvalue of $S_{0}$ is estimated as
@@
where, in the second inequality, we substitute
$v=(PB)^{t}(PB(PB)^{t})^{-1}\mu\in\mathcal{R}(A),$
which satisfies $PBv=\mu$. Here, the operators $(PB)^{t}$ and $PB(PB)^{t}$ are regarded as
$(PB)^{t}\colon W_{0}\to\mathcal{R}(A),\quad PB(PB)^{t}\colon W_{0}\to W_{0},$
in view of (15). On the other hand, the maximum eigenvalue of $S_{0}$ is estimated as
@@
$\kappa(S_{0})\lesssim\frac{H}{h}.$
This result agrees with [11, Theorem 3.2], where the condition number of $S_{0}$ was first estimated.
@@
The only technical ingredient we require is the well-known vertex–edge lemma [45, Section 4.3], stated in Lemma 7.4. In what follows, edges are understood as relatively open subsets, i.e., without their endpoints. For a subset $K\subset\partial\Omega_{j}$, the restriction $I_{K}^{0}v_{j}\in V_{j}$ of $v_{j}\in V_{j}$ is defined by
@@
    $\|I_{x}^{0}v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)\|v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\quad v_{j}\in V_{j}.$
@@
    $\|I_{e}^{0}v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}\|v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\quad v_{j}\in V_{j}.$
@@
To estimate the maximum eigenvalue of $L_{\mathrm{FETI}}S_{0}$, Corollary 4.12 with $\bar{B}=B_{D}$ implies that it suffices to bound $(A(PB_{D})^{t}PBv,(PB_{D})^{t}PBv)$ for any $v\in\mathcal{R}(A)$. Take any $v\in\mathcal{R}(A)$ and choose $\phi\in\mathcal{N}(A)$ such that $B(v+\phi)\in W_{0}$. Then we have
@@
where the first inequality follows from the definitions of $B$ and $B_{D}$ together with the triangle inequality, and the second inequality follows from Lemma 7.4. Here we require mild geometric assumptions on the domain decomposition $\{\Omega_{j}\}_{j=1}^{J}$, namely that the numbers of vertices and edges of the subdomains are uniformly bounded and that the number of subdomains sharing a common vertex is also uniformly bounded. Hence, by Corollary 4.12, we deduce that
$\lambda_{\max}(L_{\mathrm{FETI}}S_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
In summary, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{FETI}}S_{0}$ (cf. [39, Theorem 6.15]):
$\kappa(L_{\mathrm{FETI}}S_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
We note that the unscaled Dirichlet preconditioner, namely $PBA(PB)^{t}$, was first proposed in [11] and analyzed in [28]. In the unscaled case, as discussed in [28], the minimum eigenvalue admits a weaker estimate in certain scenarios.
@@ 7.2 FETI-DP
(39) $\min_{v=(v_{j})_{j=1}^{J}\in\tilde{V}}\left\{\frac{1}{2}(\tilde{A}v,v)-(\tilde{f},v)\right\},\quad\text{subject to}\quad Bv=0,$
where the matrix $\tilde{A}$ and the vector $\tilde{f}$ are defined analogously to (32); see [18, 39] for details. Here $B$ is a full-rank matrix with entries 0 and $\pm 1$ that enforces continuity along the interior of subdomain edges, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
@@
(40) $\begin{bmatrix}\tilde{A}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}\tilde{f}\\ 0\end{bmatrix}$.
In contrast to the FETI formulation (33), in (40) the matrix $\tilde{A}$ is SPD. Therefore, by eliminating $u$ in (40), we obtain the following Schur complement system corresponding to (7):
(41) $S\lambda=d,\quad\text{where}\quad S=B\tilde{A}^{-1}B^{t},\quad d=B\tilde{A}^{-1}\tilde{f}$.
In (39), since the matrix $B$ enforces continuity only at nodal points located in the interior of subdomain edges, each nodal point is affected by $B$ only once. Hence, we have the following identity [29, equation (3.2)]:
@@
As FETI-DP involves functions that are continuous at subdomain corners, we require certain Poincaré-type inequalities associated with subdomain corners in order to analyze FETI-DP. In Lemma 7.5, which summarizes results from [21, Lemma 2] and [29, Lemma 5.1], $L^{2}$ and $H^{\frac{1}{2}}$ estimates are provided for functions vanishing at subdomain corners. In what follows, let $I_{H}$ denote the linear coarse interpolation operator associated with the coarse triangulation $\mathcal{T}_{H}$.
@@
Now, we estimate the extremal eigenvalues of $S$ given in (41). By Theorem 4.1, we obtain
@@
$\lambda_{\max}(S)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right).$
@@
$\kappa(S)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right),$
which agrees with the existing result in [21, Proposition 3].
@@
$L_{\mathrm{DP}}=\bar{B}\tilde{A}\bar{B}^{t},$
By Corollary 4.7, we have $\lambda_{\min}(L_{\mathrm{DP}}S)\geq 1$. Moreover, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, it follows that
$\begin{split}(\tilde{A}\bar{B}^{t}Bv,\bar{B}^{t}Bv)&=(\tilde{A}\bar{B}^{t}B(v-I_{H}v),\bar{B}^{t}B(v-I_{H}v))\\ &\eqsim\sum_{j=1}^{J}|\bar{B}^{t}B(v-I_{H}v)|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\sum_{j=1}^{J}\sum_{e\text{:~edge of }\Omega_{j}}\left|I_{e}^{0}(v_{j}-I_{H}v_{j})\right|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\left(1+\log\frac{H}{h}\right)^{2}\sum_{j=1}^{J}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\eqsim\left(1+\log\frac{H}{h}\right)^{2}(\tilde{A}v,v),\end{split}$
@@
$\lambda_{\max}(L_{\mathrm{DP}}S)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
Finally, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{DP}}S$ (cf. [29, Theorem 5.2]):
$\kappa(L_{\mathrm{DP}}S)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
@@
Two closely related nonoverlapping domain decomposition methods to FETI and FETI-DP are BDD (balancing domain decomposition) [14, 25] and BDDC (balancing domain decomposition by constraints) [9, 14]. It has been shown in the literature [7, 24, 26, 27, 38] that FETI/FETI-DP with scaled Dirichlet preconditioners, and BDD/BDDC are spectrally equivalent in the sense that their spectra agree except for the eigenvalues 0 and $1$. In fact, BDD and BDDC can be analyzed within our framework (Theorems 4.5 and 4.10) by essentially the same arguments used for FETI and FETI-DP. For brevity, we omit the details.